OpenStudy (anonymous):

A flat rectangular piece of aluminum has a perimeter of 66 inches. The length is 15 inches longer than the width. Find the width. Question 6 options: 33 inches 9 inches 24 inches 39 inches

OpenStudy (kymber):

There's probably a better way to do this, but I think it's easy just to do a guess and check kind of thing. So, if the length is fifteen inches longer than the width, and it gives us four options for the width, then we can add fifteen to each width to find four possible length options. 1) 33 + 15 = 48 2) 9 + 15 = 24 3) 24 + 15 = 39 4) 39 + 15 = 54 Then we can find possible perimeters by adding the length + length + width + width 1) 33 + 33 + 48 + 48 = ? 2) 9 + 9 + 24 + 24 = ? 3) 25 + 24 + 39 + 30 = ? 4) 39 + 39 + 54 + 54 = ? Which one adds up to 66 inches?

OpenStudy (anonymous):

Number 2 Thanks!!!

OpenStudy (kymber):

You're welcome :D